OpenStudy (anonymous):

Suppose that the functions u and w u(x)=x^2+5 w(x)=sqrtx+7 Find the following. (u o w)(2) (w o u)(2)

OpenStudy (anonymous):

what is \(w(2)\) ?

OpenStudy (anonymous):

@ganeshie8 can you help with this?

ganeshie8 (ganeshie8):

\(\large (u~o~w)(2)\) is same as \(\large u( ~w(2)~)\)

ganeshie8 (ganeshie8):

start by finding \(w(2)\)

ganeshie8 (ganeshie8):

\(\large w(x)=\sqrt{x+7}\) \(\large w(2) = ?\)

OpenStudy (anonymous):

are you saying to substitute x for 2 then evaluate?

ganeshie8 (ganeshie8):

yup !

OpenStudy (anonymous):

\[w(x)=\sqrt{2+7}\]

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

\[=\sqrt{9}=3\]

ganeshie8 (ganeshie8):

yes !

ganeshie8 (ganeshie8):

\(\large (u~o~w)(2) = u(~w(2)~)\) \(\large ~~~~~~~~~ ~~~~~~~= u(~3~)\) \(\large ~~~~~~~~~~~~~~~~~ = ?\)

OpenStudy (anonymous):

so now I am substituting 3 for x?

ganeshie8 (ganeshie8):

yes, u do that in u(x) ok ?

OpenStudy (anonymous):

14?

ganeshie8 (ganeshie8):

\(\large (u~o~w)(2) = u(~w(2)~)\) \(\large ~~~~~~~~~ = u(~3~)\) \(\large ~~~~~~~~~ = 3^2+5 = 14\)

ganeshie8 (ganeshie8):

14 is right !

ganeshie8 (ganeshie8):

see if u can find the second one : (w o u)(2)

OpenStudy (anonymous):

4?

ganeshie8 (ganeshie8):

You got it !

ganeshie8 (ganeshie8):

good job :)